Grade 11Algebra

log 120 base 5 + (x - 3) - 2log(1 - 5^ x-3 base 5) = - log(0.2 - 5^x-4 base 5)

cbrajender
10 Years agoGrade 11

1 Answer

arvind
10 Years ago
On solving using identities and substituting 5^x-3 as 'u',we get a quadratic equation 
25t^2 -26t+1=0 which gives t=1/25 or t=1.
Resubstitute to give 5^x-3=1/25 which gives x=1
also 5^x-3=1 which gives x=3
Hope that helped!any doubt feel free to ask.
Cheers!